# Charmonium spectroscopy with optimal distillation profiles

### 27 Dec 2023 v1

@@ 1 Introduction
The method of optimal distillation profiles introduced in [1] has been used to define meson operators with large overlaps onto energy eigenstates of interest, significantly improving on standard distillation [2], and to study glueball-charmonium mixing [3] and static potentials [4]. The goal of this work is to use it to study the charmonium spectrum and its mixing with glueballs and light mesons in $N_{f}=3+1$ gauge ensembles. The significant suppression of excited-state contamination at small time separations yielded by the optimal profiles is expected to improve the signal of quark-disconnected contributions, which are necessary to study iso-scalar meson operators yet heavily affected by a signal-to-noise problem. Two different pion masses $\left(\frac{m_{\pi}}{m_{\eta_{c}}}\approx 3,7\right)$ are used, which change the decay channels for glueballs and two-particle states. A first step to study these decay dynamics is also done in this work by mapping out the energy spectrum based on single-particle operators.
@@ 2 Methods
Two ensembles of $N_{f}=3+1$ clover improved Wilson fermions, Lüscher-Weisz gauge action, open boundary conditions in time at the $SU(3)$ light flavor symmetric point are used in this work [5]. The first one, denoted as B, has size $48^{3}\times 144$, $\beta=3.43$, lattice spacing $a\approx 0.043$ fm and $m_{\pi}\approx 420$ MeV. The second one, denoted as A1-heavy, has size $32^{3}\times 96$, $\beta=3.24$, $\frac{m_{\pi}}{m_{\eta_{c}}}=3$ and lattice spacing $a\approx 0.068$ fm determined from the spin-singlet splitting $m_{h_{c}}-m_{\eta_{c}}$. The latter ensemble is particularly useful for meson-glueball mixing since the two-pion decay threshold is significantly raised while the former ensemble is particularly useful for mapping the charmonium spectrum thanks to the light quark masses being tuned to their sum in nature [6]. The observable of interest for these ensembles is the temporal correlation matrix

### 14 Jan 2024 v2

@@ 1 Introduction
The method of optimal distillation profiles introduced in [1] has been used to define meson operators with large overlaps onto energy eigenstates of interest, significantly improving on standard distillation [2], and to study glueball-charmonium mixing [3] and static potentials [4]. The goal of this work is to use it to study the charmonium spectrum and its mixing with glueballs and light mesons in $N_{f}=3+1$ gauge ensembles. The significant suppression of excited-state contamination at small time separations yielded by the optimal profiles is expected to improve the signal of quark-disconnected contributions, which are necessary to study iso-scalar meson operators yet heavily affected by a signal-to-noise problem. Two different pion masses $\left(\frac{m_{\eta_{c}}}{m_{\pi}}\approx 3,7\right)$ are used, which change the decay channels for glueballs and two-particle states. A first step to study these decay dynamics is also done in this work by mapping out the energy spectrum based on single-particle operators.
@@ 2 Methods
Two ensembles of $N_{f}=3+1$ clover improved Wilson fermions, Lüscher-Weisz gauge action, open boundary conditions in time at the $SU(3)$ light flavor symmetric point are used in this work [5, 6]. The first one, denoted as B, has size $48^{3}\times 144$, $\beta=3.43$, lattice spacing $a\approx 0.043$ fm and $m_{\pi}\approx 420$ MeV. The second one, denoted as A1-heavy, has size $32^{3}\times 96$, $\beta=3.24$, $\frac{m_{\eta_{c}}}{m_{\pi}}=3$ and lattice spacing $a\approx 0.068$ fm determined from the spin-singlet splitting $m_{h_{c}}-m_{\eta_{c}}$. The latter ensemble is particularly useful for meson-glueball mixing since the two-pion decay threshold is significantly raised while the former ensemble is particularly useful for mapping the charmonium spectrum thanks to the light quark masses being tuned to their sum in nature [7]. The observable of interest for these ensembles is the temporal correlation matrix